OpenStudy (anonymous):

What is the x value of the solution to the system of equation? Y= 4x -4 6x+ y=1

OpenStudy (shadowlegendx):

@jim_thompson5910 You got this one? ;p

jimthompson5910 (jim_thompson5910):

6x+ y=1 6x+ 4x-4=1 ... replace y with 4x-4 (since y = 4x-4) now solve for x

jimthompson5910 (jim_thompson5910):

you can go ahead if you want

OpenStudy (shadowlegendx):

@jim_thompson5910 is correct. You input in 4x - 4 for y, as y = 4x - 4 You get: 6x+ 4x-4=1 Like terms 10x - 4 = 1 Add 4 to both sides 10x = 5 Divide both sides by 10 x = 0.5

OpenStudy (shadowlegendx):

@megantownsend Link me if you need anymore help :)

OpenStudy (anonymous):

Okay thank u

OpenStudy (shadowlegendx):

No problem :)

OpenStudy (anonymous):

Wait how do u link someone?

jimthompson5910 (jim_thompson5910):

he meant tag him

jimthompson5910 (jim_thompson5910):

you type @ in front of the person's username

OpenStudy (anonymous):

Oh okay thank you!

jimthompson5910 (jim_thompson5910):

np